Gary’s TV had the following accounts and amounts in its financial statements on December 31, 2019. Assume that all balance sheet items reflect account balances at December 31, 2019, and that all income statement items reflect activities that occurred during the year then ended. Interest expense $ 5,400 Paid-in capital 17,200 Accumulated depreciation 4,800 Notes payable (long-term) 53,000 Rent expense 11,700 Merchandise inventory 124,000 Accounts receivable 41,500 Depreciation expense 2,400 Land 37,000 Retained earnings 152,500 Cash 28,000 Cost of goods sold 238,000 Equipment 28,000 Income tax expense 57,000 Accounts payable 31,000 Net sales 400,000 Required: Calculate the difference between current assets and current liabilities for Gary’s TV at December 31, 2019. Calculate the total assets at December 31, 2019. Calculate the earnings from operations (operating income) for the year ended December 31, 2019. Calculate the net income (or loss) for the year ended December 31, 2019. What was the average income tax rate for Gary’s TV for 2019? If $27,500 of dividends had been declared and paid during the year, what was the January 1, 2019, balance of retained earnings?


a.Difference

b.Total assets

c.Operating income

d. Net income or net loss

e.Average income tax rate%

f.Retained earnings
Answer #1

a. Difference = Current assets - Current liabilities

Difference = (Cash+Accounts receivable+Merchandise inventory) - Accounts payable

Difference = $28,000+41,500+124,000 - 31,000 = $162,500

b. Total assets = Cash+Accounts receivable+Merchandise inventory+Land+Equipment-accumulated depreciation

. Total assets = $28,000+41,500+124,000+37,000+28,000-4,800 = $253,700

c. Operating income :

Net sales $400,000
Cost of goods sold 238,000
Gross profit 162,000
Less: Operating expenses:
Depreciation 2,400
Rent expense 11,700
Total operating expenses 14,100
Operating income $147,900

d. Net income:

Operating income $147,900
Interest expense 5,400
Income before tax 142,500
Income tax expense 57,000
Net income $85,500

e. Average income tax rate = $57,000/142,500*100 = 40%

f. Retained earnings, January 1 = Retained earnings, December 31 + Dividends - Net income

Retained earnings, January 1 = $152,500 + 27,500 - 85,500 = $94,500